I am unsure how to solve this type of eqations 4.2 +2.5a=9.2 can you please walk me through this.thank you.
Answer by TimothyLamb(4379)

---
4.2 + 2.5a = 9.2
2.5a = 9.2 - 4.2
a = (9.2 - 4.2)/2.5
a = (5.0)/2.5
a = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php